OpenStudy (anonymous):

Mrs. Jones buys two toys for her son. The probability that the first toy is defective is 1/3, and the probability that the second toy is defective given that the first toy is defective is 1/5. What is the probability that both toys are defective?

OpenStudy (perl):

P( A & B) = P( A) * P ( B |A )

OpenStudy (anonymous):

That's the one thing I am not understanding.

OpenStudy (anonymous):

I don't understand the product rule functions.

OpenStudy (anonymous):

Or at least the placement of the probabilities.

OpenStudy (anonymous):

@perl

OpenStudy (perl):

Let A = the event that first toy is defective B = the event that second toy chosen is defective so we are given P(A) = 1/3 P(B|A) = 1/5 B|A means "event B given that you know event A has occured" now we are asked to find P( A & B ) the rule is P( A & B ) = P(A) * P(B|A) so we plug in P(A& B ) = 1/3 * 1/5